# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ 4. The CM case and magnetic modular forms
If one only focuses on $a_{p}=a_{p}(G_{k,D}^{(r)})$, then generically one obtains the following table from 4.1 and 4.2.
@@
To some extent, this also explains the $(r^{\prime}-1)$-magnetic property of $G_{k,D}^{(r)}$.
Before proceeding further with numerical phenomena, we present an example for $k=4$ and D=7D=-7, which may help with understanding the setting.
Example 4.6.
Let $k=4$ and D=7D=-7. Explicitly, we may choose $\alpha_{D}=\frac{1+\sqrt{-7}}{2}$ and $C/\mathbb{Q}$ given by $y^{2}+xy=x^{3}-x^{2}-2x-1$ with LMFDB label 49.a4 [LMF25, Elliptic Curve 49.a4]. In this case, $j(C)=j(\alpha_{D})=-3375$. Then,
$G_{4,-7}^{(1)}$ $\;=\;\frac{E_{4}}{j+3375}$
$G_{4,-7}^{(2)}$ = $\;=\;19\cdot\frac{E_{4}}{j+3375}-91125\cdot\frac{E_{4}}{(j+3375)^{2}}$
$G_{4,-7}^{(3)}$ = $\;=\;1399\cdot\frac{E_{4}}{j+3375}-19008675\cdot\frac{E_{4}}{(j+3375)^{2}}+54251268750\cdot\frac{E_{4}}{(j+3375)^{3}}.$
For all primes $p$ with $(\frac{-7}{p})=1$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{4,-7}^{(1)})$ $\;\equiv\;u_{p}(C)^{2}{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(1)})$ $\pmod{p^{3l}}$
$a_{np^{l}}(G_{4,-7}^{(2)})$ $\;\equiv\;p{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(2)})$ $\pmod{p^{3l}}$
$a_{np^{l}}(G_{4,-7}^{(3)})$ $\;\equiv\;u_{p}(C)^{-2}p^{2}{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(3)})$ $\pmod{p^{3l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit. Moreover, $G_{4,-7}^{(2)}$ is $1$-magnetic.
In fact, one can already observe interesting numerical phenomena for the family $\{G_{k,D}^{(r)}\}_{1\leq r\leq k-1}$ by viewing $\mathrm{Sym}^{k-2}C$ simply as a Galois representation of $\mathbb{Q}$. Indeed, we have the following splitting (as Galois representations of $\mathbb{Q}$)  [^7] An easy way of observing this splitting is through the splitting of the $L$-function of $\mathrm{Sym}^{k-2}C$.
$\mathrm{Sym}^{k-2}C\;\cong\;\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}\oplus{\hskip 1.0pt}\bigoplus_{r=1}^{\frac{k-2}{2}}\big(\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}\big),$
where $(\frac{D}{\cdot})$ denotes (the Dirichlet character given by) the Kronecker symbol and $\varepsilon_{\ell}$ denotes the $\ell$-adic cyclotomic character. Transferring this splitting to the modular form side, one also obtains certain congruences for $G_{k,D}^{(k/2)}$ and the others. In this case, it is more convenient to write everything in terms of negative discriminants.
@@
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Then, $G_{k,D}^{(k/2)}$ satisfies that for all primes $p$ with $p\nmid D$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(k/2)})\;\equiv\;\big(\big(\tfrac{D}{p}\big){\hskip 1.0pt}p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(G_{k,D}^{(k/2)})\pmod{p^{(k-1)l}}$
@@
We will prove this in Section 6. Note that there are no assumptions on $v_{p}(j(C))$ and $v_{p}(j(C)-1728)$ so primes dividing $D$ are the only “bad primes”. In fact, we will also prove some $p$-adic property for $p{\hskip 1.0pt}|{\hskip 1.0pt}D$, from which the $\frac{k-2}{2}$-magnetic property of (some multiple of) $G_{k,D}^{(k/2)}$ as predicted by 4.4 then follows.
@@
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$ with $r\neq\frac{k}{2}$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form (in fact Hecke eigenform)
@@
This should be viewed as a natural generalization of the ASD congruence observed in the case of CM elliptic curves and simple poles, i.e., 2.6 or 2.8. As in Remark 2.7, 4.9 also follows from the supersingular (4.1) and ordinary (4.2) dichotomy.
The final observations we have on the family $\{G_{k,D}^{(r)}\}_{1\leq r\leq k-1}$ remain somewhat mysterious to us. In the splitting (7) of the elliptic curve side over $K$, $\psi_{C}^{a}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}\overline{\psi_{C}^{a}}$ multiply to $(\psi_{C}\overline{\psi_{C}})^{k-2}=(\mathrm{Res}_{K/\mathbb{Q}}\varepsilon_{\ell})^{k-2}$. It seems that this is also reflected on the modular form side via the following phenomena.
@@
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l}}$.
@@
It is easy to check that the case when $r=k/2$ follows from Theorem 4.7 and the case when $(\frac{D}{p})=1$ follows from 4.2. When $(\frac{D}{p})=-1$, the numerical observations also seem to suggest that for all $n,l\in\mathbb{Z}^{+}$ with $l\geq 2$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l+(k-2r^{\prime})}}$.

### v2

@@ 4.2. The magnetic property and other observations
If one only focuses on $a_{p}=a_{p}(G_{k,D}^{(r)})$, then generically one obtains the following table from 3.4 (or 4.5) and 4.1.
@@
To some extent, this provides a visual explanation of the $(r^{\prime}-1)$-magnetic property of $G_{k,D}^{(r)}$.
@@
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, $G_{k,D}^{(k/2)}$ satisfies that for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(k/2)})\;\equiv\;\big(\big(\tfrac{D}{p}\big){\hskip 1.0pt}p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(G_{k,D}^{(k/2)})\pmod{p^{(k-1)l}}$.
@@
We include this result here for the completeness of the discussion. It simply follows from Theorem 1.10, since $G_{k,D}^{(k/2)}$ is a nonzero rational multiple of $\widetilde{\mathcal{G}}_{k,D}$, and the rational multiple involves only divisors of $D$, $j(\alpha_{D})$, and $j(\alpha_{D})-1728$.
@@
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$ with $r\neq\frac{k}{2}$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form
@@
This should be viewed as a generalization of the three-term ASD congruence observed in the case of CM elliptic curves and simple poles, i.e., 2.4 or 2.6. As in Remark 2.5, this conjecture also follows from the supersingular (3.4) and ordinary (4.1) dichotomy.
The last two observations are related to tensor products. In the splitting (8) of the elliptic curve side over $K$, one has that $\psi_{C}^{a}\overline{\psi_{C}^{b}}\otimes\psi_{C}^{b}\overline{\psi_{C}^{a}}=(\psi_{C}\overline{\psi_{C}})^{k-2}=(\mathrm{Res}_{K/\mathbb{Q}}\varepsilon_{\ell})^{k-2}$. On the modular form side, this suggests that the “tensor product” of $G_{k,D}^{(r)}$ and $G_{k,D}^{(k-r)}$ should resemble the character $\varepsilon_{\ell}^{k-2}$. Here we form the “tensor product” of two modular forms via termwise multiplication, motivated by the Rankin–Selberg theory.
@@
Let $k\in\{4,6,8,10,14\}$, $r\in\mathbb{Z}$ with $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l}}$. (15)
@@
When generalizing this conjecture to other weights in Section 5, it is possible that in the case when $(\frac{D}{p})=-1$, one may need to multiply both sides by $(r-1)!\cdot(k-1-r)!$ to account for the factors that appear in 4.5. Also, when $(\frac{D}{p})=-1$, the numerical observations suggest that the congruence (15) holds modulo $p^{(k+r^{\prime}-2)l+(k-2r^{\prime})}$ for $l\geq 2$.